# Computing an Aircraft’s Gliding Range and Minimal Return Altitude in Presence of Obstacles and Wind

### v1

@@ A.2 Derivaton of HJB equations for MRAP
To better understand this derivation we advise the reader to first go through section II.2. In this section we rename $V_{3}\to V$ in order to simplify the notation. Let $\bm{x}:[0,h_{f}]\mapsto\operatorname{\mathbb{R}}^{2}$ be the path joining $\bm{x}(0)=\bm{x}_{a}$ and $\bm{x}(h_{f})=\bm{y}$ that achieves the minimum in the problem (30). Then we can apply Bellman’s optimality principle. In the following we name $\bm{\nu}\coloneqq\frac{d\bm{x}}{dh}(h_{f})$. Notice $\bm{\nu}\in\operatorname{\mathbb{R}}^{2}$.
@@
In the last step the minimization is subject to the constraints
@@
The by case solution is needed to appropriately enforce that the aircraft does not fall below $h_{\text{min}}$. The second case is the simplest. If the aircraft is above $h_{\text{min}}$ then it shall continue climbing with steepness $1/\tilde{g}$. If instead the aircraft is at $h_{\text{min}}$, then continuing the climb with a steepness of $1/\tilde{g}$ might result in falling below $h_{\text{min}}$ in case the terrain rises more steeply. When going in direction $\bm{\hat{a}}$ the terrain steepness is given by $\nabla h_{\text{min}}(\bm{y})\cdot\bm{\hat{a}}$. Since the steepness of the path is $1/\left\lVert d\bm{x}/dh\right\rVert=1/\left\lVert\bm{\nu}\right\rVert$, we have that $1/\left\lVert\bm{\nu}\right\rVert\geq\max\left(1/\tilde{g},\nabla h_{\text{min}}(\bm{y})\cdot\bm{\hat{a}}\right)$.
@@ A.3 Simplifications of the HJB equations for MRAP
In this section we aim to derive simplified forms of (43), with the goal of better understanding it. We start from the one dimensional case, without wind. When $V>h_{\text{min}}$ the equation is $|dV/dy(y)|>1/g$. We now consider the case $V=h_{\text{min}}$.
@@
Figure 6 shows the behavior of $V$ in a one dimensional setting. From the figure we observe that $V$ grows with steepness $1/g$ until it gets to the same height of $h_{\text{min}}$. Then it starts tracking $h_{\text{min}}$ as long as $dh_{\text{min}}(y)/dy>1/g$. This captures a central feature of MRAP. In absence of obstacles and with constant wind, $V$ forms a cone around $\bm{x}_{a}$ with steepness $1/g$. From this figure we notice an important feature of the solution: even if $h_{\text{min}}$ is differentiable, $V$ is not. In the points where we transititon from the cone solution to tracking the terrain the derivative of $V$ is discontinuous.
@@
We shall now consider the 2d case in absence of wind. Looking at equation (43) let’s first consider the case $V(\bm{y})>h_{\text{min}}(\bm{y})$ (no interaction with obstacles). In this case the aircraft climbs with steepness $1/g$. In absence of wind, $g$ is constant with respect to $\bm{\hat{a}}$ and $V(\bm{y})=h_{\text{min}}(\bm{x}_{a})+\frac{1}{g}\left\lVert\bm{y}-\bm{x}_{a}\right\rVert$.
@@ Appendix B Analytical solution of GRRP and MRAP
In the case of uniform wind and no elevation (i.e. $h_{\text{min}}(\bm{x})=-\infty\;\forall\bm{x}$ ) we can solve both the GRRP and the MRAP equations exactly. We show this in the case of GRRP, but the solution for MRAP is identical. Suppose $\bm{W}$ is the wind vector in the whole space. Let $z_{0}\in\operatorname{\mathbb{R}}$ be the initial altitude of the glider, and $\bm{x}_{0}$ its initial position. Since the wind is uniform, the function $g$ depends only on $\bm{\hat{a}}$. Equations (13) and (14) then become
@@ Appendix C Algorithms for GRRP
In this section we present an algorithm to solve GRRP in absence of wind. This pseudocode does not reflect the actual implementation of Glikonal-G in the github repository. In particular Algorithm 1 is based on the FMM while the repository implementation is based on OUMs. We still find it valuable to include this pseudocode as it shows how the algorithm works in a simplified setting, namely the case of zero wind. The OUM based implementation can account for an arbitrary wind vector field, but it is also more complex. The computational complexity of both 1 and its OUM version is $O(n\log n)$ where $n$ is the number of nodes in the graph.
@@
In this algorithm the 2d space is discretized into a grid with $n$ nodes. $h$ is the spacing between adjacent nodes on the grid. The edges in the grid are represented through an adjacency list ’$N$’. The list $E$ contains the function $h_{\text{min}}$ evaluated on the grid sites.
The algorithms works by dividing the nodes into three classes: far, considered and accepted. Far nodes are those that have not been yet reached by the expanding front. Accepted ones are those for which $U$ is known. Considered nodes are nodes on the front. Initially only the node corresponding to the initial position of the glider is in the considered set. At each step in the while loop we take the considered node with the lowest value of $U$ and move it into the accepted nodes. Then we compute a tentative value $\tilde{U}$ for its neighbours. If $z_{0}-\tilde{U}>h_{\text{min}}$ then the neighbour is added to the considered nodes (if it wasn’t already among the considered nodes). The process halts when there are no more considered nodes.
To find the minimum in $O(\log n)$ time in line 10, a heap should be used to store the considered nodes. In the case of no wind, the vector field $\bm{\hat{a}}^{\star}_{G}$, defined in (15) is equal to $\nabla U$. Below, the implementation of the function EikonalUpdate
@@
8:     $\tilde{U}\leftarrow\min\left(U_{x},U_{y}\right)+h/g$triangleright When the quadratic does not admit a solution use this one dimensional update
@@ Appendix D Altitude loss in turns
The modeling of optimal control problem in both GRRP and MRAP presents some differences from the real behavior of a real aircraft.
@@ D.1 Modeling the altitude loss in a turn
Many works [6, 7] that aim to compute the gliding range in absence of obstacles assume that the glider will take an initial turn at constant bank angle from the initial heading towards the final heading. Once the turn is completed the glider continues in a straight line.
@@
In this case the glider’s initial position is $(0,0)$ and its heading is indicated by the black arrow. $\bm{x}$ is the final position. The red line represents the optimal trajectory under our model of the dynamics. The concatenation of the blue and green lines is instead a more reasonable trajectory that can actually be flown.
@@
Our goal is to compare the altitude loss on the red trajectory with that of the green-blue one, to understand in which cases the initial turn contributes significantly to the total altitude loss. We assume that the glider’s trajectory is composed by an initial circular arc, of fixed radius $R$, followed by a straight line. We indicate with $g_{R}$ the glide ratio during the turn, and with $g_{\infty}$ the glide ratio during straight flight. indicating with $\Delta H_{t},\Delta H$ respectively the altitude losses on the green-blue and red trajectories.
@@
At this point it is natural to wonder if there is a way to compensate for this difference in altitude loss. The simplest option is to use a more conservative glide ratio (i.e., a smaller one) for the computation of $\Delta H$. Naming this $\tilde{g}_{\infty}<g_{\infty}$, we have $\Delta H_{t}-\Delta H=\frac{\left\lVert\bm{r}_{2}\right\rVert}{g_{\infty}}-\frac{\left\lVert\bm{x}\right\rVert}{\tilde{g}_{\infty}}+R\phi/g_{R}$. When $\left\lVert\bm{x}\right\rVert$ is sufficiently large, this quantity is negative. This option however is not satisfactory when $\left\lVert\bm{x}\right\rVert$ is too small.
To overcome this problem we can use a different approach. As we presented it, the Glikonal-G algorithm starts expanding the front from a single node. In general however Glikonal-G can also take as input a partial solution, i.e., a function $U$ whose values are defined only on a certain region of space. Then Glikonal-G expands the solution to the whole space.
@@ Appendix E Algorithm for MRAP
In this appendix we provide the pseudocode of the Glikonal-M algorithm. Glikonal-M works on a grid with identical horizontal and vertical spacings. This implementation is does not take wind into account and solves equation (48). It is based on the Fast Marching Method. The following pseudocode reflects the algorithm published in the repository.
Algorithm 3 Glikonal-M. Algorithm to solve MRAP in absence of wind.
@@
In this algorithm the 2d space is discretized into a grid with $n$ nodes. $h$ is the spacing between adjacent nodes on the grid. The edges in the grid are represented through an adjacency list ’$N$’. The list $E$ contains the function $h_{\text{min}}$ evaluated on the grid sites.
The algorithms works by dividing the nodes into three classes: far, considered and accepted. Far nodes are those that have not been yet reached by the expanding front. Accepted ones are those for which $V$ is known. Considered nodes are nodes on the front. Initially only the node corresponding to the position of the airfield is in the considered set. At each step in the while loop we take the considered node with the lowest value of $V$ and move it into the accepted nodes. Then we add its neighbours to the considered set (if they were not there already) and compute a tentative value of $V$ for each neighbour. The tentative value is computed taking the maximum between the elevation $E$ and the altitude needed to glide back to the airport in absence of the obstacle. The process halts when there are no more considered nodes.
@@ Appendix F Additional numerical experiments
In this appendix we present an additional experimental setting for Glikonal-G and Glikonal-M. We consider specifically the case of an elevation profile including a mountain range with two passes. In both cases the glide ratio in absence of wind is one and the grid has size $101\times 101$, with a grid spacing of one.
@@ F.1 GRRP
Figure 8: Glikonal-G’s solution in the case of non uniform wind and a mountain range with two passes. The red dot represents the glider’s initial position. (a) Wind direction (constant with altitude) plotted on top of the elevation profile (displayed as a heatmap). (b) Contour lines of the function $z_{0}-U_{G}$ (in white) and optimal trajectories (in turquoise). The heatmap is the elevation profile, and the red shaded regions are those not reachable in gliding flight. (c) 3d representation of the function $z_{0}-U_{G}$, in orange and of $h_{\text{min}}$.
@@ F.2 MRAP
We test Glikonal-G on an elevation profile depicted in figure 9(a) as a heatmap. The purple dot indicates the position of the airfield. In the same plot, the contour lines of $V$ are shown in white and the feasible re-entry paths in turquoise. We remark that the re-entry trajectories tend to pass through one of the two saddles. In a small region just behind the middle of the mountain range, it is then convenient to pass the range directly, since reaching a saddle would lead to a greater loss of altitude. This region is well visible in panel (b), where it takes a triangular shape, with the base at the top of the mountain range.
@@ Appendix G Error in Glikonal-M and Glikonal-G
In this appendix we test the Glikonal-M and Glikonal-G algorithms in several settings where we can compute the true solution. The errors are due to the discretization of the PDE, which is numerically solved on a grid. We also verify that when the grid spacing decreases the error also does. For proofs of convergence of the discretized solution to the true one when the grid spacing approaches zero we refer the reader to [1].
@@ G.1 Glikonal-G
Name $U_{G}$ the approximation of $U$ outputted by Glikonal-G. The following experiments are all run on a $101\times 101$ grid with a spacing of one. The glide ratio is also one in absence of wind. When wind is present the glide ratio in direction $\bm{\hat{a}}$ is $g(\bm{x},y,\bm{\hat{a}})=\left\lVert\bm{\hat{a}}+\bm{W}(\bm{x},z)\right\rVert$, corresponding to the fact that the glider moves in every direction with the same airspeed of one. The red dot indicates the initial position of the glider.
@@
In the second experiment we place an infinitely high barrier in the middle of the domain. We leave two small openings where $h_{\text{min}}=0$ respectively in the upper and lower portion of the barrier. Apart from the barrier we set $h_{\text{min}}=0$ everywhere else. The barrier is depicted as a white vertical line in figure 11. The initial altitude is 120. The wind is uniform on strength $0.4$ and pointing north. In (a) we plot $z_{0}-U_{G}$. The panels (b,c) instead show the relative and total error, defined as above. The total error reaches its maximum value on the boundary of the GRR. In figure 16 we give a numerical demonstration of the fact that when the grid spacing goes to zero, $U_{G}$ approaches $U$. We used a grid with size $404\times 404$ and spacing 0.25. By comparing panels (b,c) with those of the previous figure we observe that both the relative and total error are greatly reduced. On the boundary of the GRR some small errors remain.
@@
Figure 12: Test of Glikonal-G on elevation profile with an infinite barrier with two passes, same as in figure 11. The grid spacing is 0.25, one quarter of the grid spacing in the previous figure and the grid size is $404\times 404$. (a) Heatmap and contour lines of the function $z_{0}-U_{G}$. The turquoise lines are the optimal trajectories and the red dot is the initial position of the glider. (b) Relative error, $(U_{G}-U)/U$. (c) Total error $U_{G}-U$.
In the last test we consider we place a barrier in the same position as the previous experiment. This time however the barrier has a finite elevation of 45. This way the glider can pass the barrier if it approaches it directly, but cannot do so if it proceeds diagonally. The two white segments represent where the barrier stops the glider from passing. In the rest of space we set $h_{\text{min}}=0$. The wind has a strength of 0.3 and is directed towards the north. In panel (a) the function $z_{0}-U_{G}$ is plotted. The panels (b,c) instead show the relative and total error, defined as above. The maximum error is attained on the bottom right. In this case the error arises from the fact that in the true solution the glider manages to pass the barrier in a slightly larger (by one pixel) portion of it.
Figure 13: Test of Glikonal-G on elevation profile with an finite barrier of height 45. The glider has enough altitude to traverse the barrier in the middle but not on the sides. The two white segments represent where the glider fails to cross the barrier. (a) Heatmap and contour lines of the function $z_{0}-U_{G}$. The turquoise lines are the optimal trajectories and the red dot is the initial position of the glider. (b) Relative error, $(U_{G}-U)/U$. (c) Total error $U_{G}-U$.
In all experiments the relative error of Glikonal-G is below 4%. Moreover all errors overestimate the altitude loss, leading to conservative estimates of the GRR.

### v2

@@ A.2 Derivaton of HJB equations for MRAP
To better understand this derivation we advise the reader to first go through section II.2. In this section, we rename $V_{3}\to V$ in order to simplify the notation. Let $\bm{x}:[0,h_{f}]\mapsto\operatorname{\mathbb{R}}^{2}$ be the path joining $\bm{x}(0)=\bm{x}_{a}$ and $\bm{x}(h_{f})=\bm{y}$ that achieves the minimum in the problem (30). Then we can apply Bellman’s optimality principle. In the following we name $\bm{\nu}\coloneqq\frac{d\bm{x}}{dh}(h_{f})$. Notice $\bm{\nu}\in\operatorname{\mathbb{R}}^{2}$.
@@
In the last step, the minimization is subject to the constraints
@@
The by-case solution is needed to appropriately enforce that the aircraft does not fall below $h_{\text{min}}$. The second case is the simplest. If the aircraft is above $h_{\text{min}}$ then it shall continue climbing with steepness $1/\tilde{g}$. If instead the aircraft is at $h_{\text{min}}$, then continuing the climb with a steepness of $1/\tilde{g}$ might result in falling below $h_{\text{min}}$ in case the terrain rises more steeply. When going in direction $\bm{\hat{a}}$ the terrain steepness is given by $\nabla h_{\text{min}}(\bm{y})\cdot\bm{\hat{a}}$. Since the steepness of the path is $1/\left\lVert d\bm{x}/dh\right\rVert=1/\left\lVert\bm{\nu}\right\rVert$, we have that $1/\left\lVert\bm{\nu}\right\rVert\geq\max\left(1/\tilde{g},\nabla h_{\text{min}}(\bm{y})\cdot\bm{\hat{a}}\right)$.
@@ A.3 Simplifications of the HJB equations for MRAP
In this section, we aim to derive simplified forms of (43), with the goal of better understanding it. We start from the one dimensional case, without wind. When $V>h_{\text{min}}$ the equation is $|dV/dy(y)|>1/g$. We now consider the case $V=h_{\text{min}}$.
@@
Figure 6 shows the behavior of $V$ in a one-dimensional setting. From the figure we observe that $V$ grows with steepness $1/g$ until it gets to the same height of $h_{\text{min}}$. Then it starts tracking $h_{\text{min}}$ as long as $dh_{\text{min}}(y)/dy>1/g$. This captures a central feature of MRAP. In the absence of obstacles and with constant wind, $V$ forms a cone around $\bm{x}_{a}$ with steepness $1/g$. From this figure, we notice an important feature of the solution: even if $h_{\text{min}}$ is differentiable, $V$ is not. In the points where we transition from the cone solution to tracking the terrain the derivative of $V$ is discontinuous.
@@
We shall now consider the 2d case in the absence of wind. Looking at equation (43) let’s first consider the case $V(\bm{y})>h_{\text{min}}(\bm{y})$ (no interaction with obstacles). In this case, the aircraft climbs with steepness $1/g$. In the absence of wind, $g$ is constant with respect to $\bm{\hat{a}}$ and $V(\bm{y})=h_{\text{min}}(\bm{x}_{a})+\frac{1}{g}\left\lVert\bm{y}-\bm{x}_{a}\right\rVert$.
@@ Appendix B Analytical solution of GRRP and MRAP
In the case of uniform wind and no elevation (i.e. $h_{\text{min}}(\bm{x})=-\infty\;\forall\bm{x}$ ) we can solve both the GRRP and the MRAP equations exactly. We show this in the case of GRRP, but the solution for MRAP is identical. Suppose $\bm{W}$ is the wind vector in the whole space. Let $z_{0}\in\operatorname{\mathbb{R}}$ be the initial altitude of the glider and $\bm{x}_{0}$ its initial position. Since the wind is uniform, the function $g$ depends only on $\bm{\hat{a}}$. Equations (13) and (14) then become
@@ Appendix C Algorithms for GRRP
In this section, we present an algorithm to solve GRRP in the absence of wind. This pseudocode does not reflect the actual implementation of Glikonal-G in the GitHub repository. In particular Algorithm 1 is based on the FMM while the repository implementation is based on OUMs. We still find it valuable to include this pseudocode as it shows how the algorithm works in a simplified setting, namely the case of zero wind. The OUM based implementation can account for an arbitrary wind vector field, but it is also more complex. The computational complexity of both 1 and its OUM version is $O(n\log n)$ where $n$ is the number of nodes in the graph.
@@
In this algorithm, the 2d space is discretized into a grid with $n$ nodes. $h$ is the spacing between adjacent nodes on the grid. The edges in the grid are represented through an adjacency list ’$N$’. The list $E$ contains the function $h_{\text{min}}$ evaluated on the grid sites.
The algorithms works by dividing the nodes into three classes: far, considered, and accepted. Far nodes are those that have not been yet reached by the expanding front. Accepted ones are those for which $U$ is known. Considered nodes are nodes on the front. Initially, only the node corresponding to the initial position of the glider is in the considered set. At each step in the while loop, we take the considered node with the lowest value of $U$ and move it into the accepted nodes. Then we compute a tentative value $\tilde{U}$ for its neighbours. If $z_{0}-\tilde{U}>h_{\text{min}}$ then the neighbour is added to the considered nodes (if it wasn’t already among the considered nodes). The process halts when there are no more considered nodes.
To find the minimum in $O(\log n)$ time in line 10, a heap should be used to store the considered nodes. In the case of no wind, the vector field $\bm{\hat{a}}^{\star}_{G}$, defined in (15) is equal to $\nabla U$. Below we give the implementation of the function EikonalUpdate
@@
8:     $\tilde{U}\leftarrow\min\left(U_{x},U_{y}\right)+h/g$triangleright When the quadratic does not admit a solution use this one-dimensional update
@@ Appendix D Altitude loss in turns
The modeling of optimal control problems in both GRRP and MRAP presents some differences from the real behavior of a real aircraft.
@@ D.1 Modeling the altitude loss in a turn
Many works [6, 7] that aim to compute the gliding range in the absence of obstacles assume that the glider will take an initial turn with constant bank angle from the initial heading towards the final heading. Once the turn is completed the glider continues in a straight line.
@@
In this case, the glider’s initial position is $(0,0)$ and its heading is indicated by the black arrow. $\bm{x}$ is the final position. The red line represents the optimal trajectory under our model of the dynamics. The concatenation of the blue and green lines is instead a more reasonable trajectory that can actually be flown.
@@
Our goal is to compare the altitude loss on the red trajectory with that of the green-blue one, to understand in which cases the initial turn contributes significantly to the total altitude loss. We assume that the glider’s trajectory is composed of an initial circular arc, of fixed radius $R$, followed by a straight line. We indicate with $g_{R}$ the glide ratio during the turn, and with $g_{\infty}$ the glide ratio during straight flight. indicating with $\Delta H_{t},\Delta H$ respectively the altitude losses on the green-blue and red trajectories.
@@
At this point, it is natural to wonder if there is a way to compensate for this difference in altitude loss. The simplest option is to use a more conservative glide ratio (i.e., a smaller one) for the computation of $\Delta H$. Naming this $\tilde{g}_{\infty}<g_{\infty}$, we have $\Delta H_{t}-\Delta H=\frac{\left\lVert\bm{r}_{2}\right\rVert}{g_{\infty}}-\frac{\left\lVert\bm{x}\right\rVert}{\tilde{g}_{\infty}}+R\phi/g_{R}$. When $\left\lVert\bm{x}\right\rVert$ is sufficiently large, this quantity is negative. This option however is not satisfactory when $\left\lVert\bm{x}\right\rVert$ is too small.
To overcome this problem we can use a different approach. As we presented it, the Glikonal-G algorithm starts expanding the front from a single node. In general, however, Glikonal-G can also take as input a partial solution, i.e., a function $U$ whose values are defined only on a certain region of space. Then Glikonal-G expands the solution to the whole space.
@@ Appendix E Algorithm for MRAP
In this appendix, we provide the pseudocode of the Glikonal-M algorithm. Glikonal-M works on a grid with identical horizontal and vertical spacings. This implementation is does not take wind into account and solves equation (48). It is based on the Fast Marching Method. The following pseudocode reflects the algorithm published in the repository.
Algorithm 3 Glikonal-M. Algorithm to solve MRAP in the absence of wind.
@@
In this algorithm, the 2d space is discretized into a grid with $n$ nodes. $h$ is the spacing between adjacent nodes on the grid. The edges in the grid are represented through an adjacency list ’$N$’. The list $E$ contains the function $h_{\text{min}}$ evaluated on the grid sites.
The algorithms works by dividing the nodes into three classes: far, considered, and accepted. Far nodes are those that have not been yet reached by the expanding front. Accepted ones are those for which $V$ is known. Considered nodes are nodes on the front. Initially, only the node corresponding to the position of the airfield is in the considered set. At each step in the while loop, we take the considered node with the lowest value of $V$ and move it into the accepted nodes. Then we add its neighbours to the considered set (if they were not there already) and compute a tentative value of $V$ for each neighbour. The tentative value is computed by taking the maximum between the elevation $E$ and the altitude needed to glide back to the airport in the absence of the obstacle. The process halts when there are no more considered nodes.
@@ Appendix F Additional numerical experiments
In this appendix, we present an additional experimental setting for Glikonal-G and Glikonal-M. We consider specifically the case of an elevation profile including a mountain range with two passes. In both cases, the glide ratio in the absence of wind is one and the grid has size $101\times 101$, with a grid spacing of one.
@@ F.1 GRRP
Figure 8: Glikonal-G’s solution in the case of non-uniform wind and a mountain range with two passes. The red dot represents the glider’s initial position. (a) Wind direction (constant with altitude) plotted on top of the elevation profile (displayed as a heatmap). (b) Contour lines of the function $z_{0}-U_{G}$ (in white) and optimal trajectories (in turquoise). The heatmap is the elevation profile, and the red-shaded regions are those not reachable in gliding flight. (c) 3d representation of the function $z_{0}-U_{G}$, in orange and of $h_{\text{min}}$.
@@ F.2 MRAP
We test Glikonal-G on an elevation profile depicted in figure 9(a) as a heatmap. The purple dot indicates the position of the airfield. In the same plot, the contour lines of $V$ are shown in white, and the feasible re-entry paths in turquoise. We remark that the re-entry trajectories tend to pass through one of the two saddles. In a small region just behind the middle of the mountain range, it is then convenient to pass the range directly, since reaching a saddle would lead to a greater loss of altitude. This region is well visible in panel (b), where it takes a triangular shape, with the base at the top of the mountain range.
@@ Appendix G Error in Glikonal-M and Glikonal-G
In this appendix, we test the Glikonal-M and Glikonal-G algorithms in several settings where we can compute the true solution. The errors are due to the discretization of the PDE, which is numerically solved on a grid. We also verify that when the grid spacing decreases the error also does. For proofs of convergence of the discretized solution to the true one when the grid spacing approaches zero we refer the reader to [1].
@@ G.1 Glikonal-G
Name $U_{G}$ the approximation of $U$ outputted by Glikonal-G. The following experiments are all run on a $101\times 101$ grid with a spacing of one. The glide ratio is also one in the absence of wind. When wind is present the glide ratio in direction $\bm{\hat{a}}$ is $g(\bm{x},y,\bm{\hat{a}})=\left\lVert\bm{\hat{a}}+\bm{W}(\bm{x},z)\right\rVert$, corresponding to the fact that the glider moves in every direction with the same airspeed of one. The red dot indicates the initial position of the glider.
@@
In the second experiment, we place an infinitely high barrier in the middle of the domain. We leave two small openings where $h_{\text{min}}=0$ respectively in the upper and lower portion of the barrier. Apart from the barrier we set $h_{\text{min}}=0$ everywhere else. The barrier is depicted as a white vertical line in figure 11. The initial altitude is 120. The wind is uniform with strength $0.4$ and pointing north. In (a) we plot $z_{0}-U_{G}$. The panels (b,c) instead show the relative and total error, defined as above. The total error reaches its maximum value on the boundary of the GRR. In figure 16 we give a numerical demonstration of the fact that when the grid spacing goes to zero, $U_{G}$ approaches $U$. We used a grid with size $404\times 404$ and spacing 0.25. By comparing panels (b,c) with those of the previous figure we observe that both the relative and total error are greatly reduced. On the boundary of the GRR some small errors remain.
@@
Figure 12: Test of Glikonal-G on elevation profile with an infinite barrier with two passes, same as in figure 11. The grid spacing is 0.25, one quarter of the grid spacing in the previous figure, and the grid size is $404\times 404$. (a) Heatmap and contour lines of the function $z_{0}-U_{G}$. The turquoise lines are the optimal trajectories and the red dot is the initial position of the glider. (b) Relative error, $(U_{G}-U)/U$. (c) Total error $U_{G}-U$.
In the last test, we place a barrier in the same position as the previous experiment. This time however the barrier has a finite elevation of 45. This way the glider can pass the barrier if it approaches it directly, but cannot do so if it proceeds diagonally. The two white segments represent where the barrier stops the glider from passing. In the rest of space, we set $h_{\text{min}}=0$. The wind has a strength of 0.3 and is directed towards the north. In panel (a) the function $z_{0}-U_{G}$ is plotted. The panels (b,c) instead show the relative and total error, defined as above. The maximum error is attained on the bottom right. In this case, the error arises from the fact that in the true solution, the glider manages to pass the barrier in a slightly larger (by one pixel) portion of it.
Figure 13: Test of Glikonal-G on elevation profile with a finite barrier of height 45. The glider has enough altitude to traverse the barrier in the middle but not on the sides. The two white segments represent where the glider fails to cross the barrier. (a) Heatmap and contour lines of the function $z_{0}-U_{G}$. The turquoise lines are the optimal trajectories and the red dot is the initial position of the glider. (b) Relative error, $(U_{G}-U)/U$. (c) Total error $U_{G}-U$.
In all experiments, the relative error of Glikonal-G is below 4%. Moreover, all errors overestimate the altitude loss, leading to conservative estimates of the GRR.